OpenStudy (anonymous):

Determine the measures of the unknown angles in the illustrations below.... It's in degrees, and that is a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

u there? lol

OpenStudy (anonymous):

65+15x-4+4x+5=180 19x=180+4-65-5 19x=114 x=114/19=6

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ur wlcum...